# Stability of Non-Hyperbolic Sets

### v1

@@ 9.3 Properties of the set AA
$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}(j)\right)=O_{2}.$
@@
Proposition 10.
Let $\ell$ be a leave of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the map $f:A(f)\mapsto A(f)$ is an homeomorphism.
@@
Since $f$ is injective, its restriction to $A$ is injective as well. We only need to prove that $f(A(f))=A(f)$. Observe that for all $x$, $f(\omega(x))=\omega(x)$. In particular, when $x=c$, $f(\omega(c))=\omega(c)$ and by Lemma 24 we obtain $f(A(f))=A(f)$. The proof is complete. ∎
@@ 10 Topological Equivalence along the Leaves
The aim of this section is to prove that two maps which are in the same leave of the lamination are conjugate on their respective Cantor sets, see Theorem 4. In order to do that, we need to build a topological model for the invariant Cantor sets.
@@ 10.1 Inverse limits
Take a leave $\ell$ of the lamination $\mathfrak{L}$ as constructed in (9.1) and (9.2). By the construction following Proposition 7 we know that for all $g$ there is a $(1,N_{g},M_{g})$-directed graph and a graph morphism $w_{g}:X_{g+1}\mapsto X_{g}$. We can then define the inverse limit as
@@
with the corresponding topology.
A sequence $\left\{\underline{x}_{n}\right\}_{n\in\mathbb{N}}=\left\{\left(x_{g,n}\right)\right\}_{n\in\mathbb{N}}\subset X$ converges to a point $\underline{x}=\left(x_{g}\right)\in X$, namely
@@
Lemma 26.
Let $\underline{x}\in X$, $\underline{x}=(x_{g})$. If there exist $g_{0}$ such that $x_{g_{0}}=q_{g_{0}}$ then for all $g\leq g_{0}$, $x_{g}=q_{g}$.
@@
In the following proposition we prove the existence of an homeomorphism on the inverse limit defined by the inverse graphs.
Proposition 11.
@@
Let $\underline{x}=(x_{g})\in X$. By Lemma 26, either $\underline{x}=\underline{q}=(q_{g})$ or there exists $g_{0}$ such that, for all $g\geq g_{0}$, $x_{g}\neq q_{g}$. For the case when $\underline{x}=\underline{q}$, define $\phi(\underline{x})=\underline{q}$. In the other case define $\phi(\underline{x})=\underline{y}=(y_{g})$ where $y_{g}=q_{g}$ for all $g<g_{0}$ and $y_{g}$ is the vertex in the unique outgoing edge starting in $x_{g}$. Since $w_{g}(x_{g+1})=x_{g}$ and since $x_{g}$ has a unique outgoing edge, if $w_{g}(\overrightarrow{x_{g+1}y_{g+1}})=\overrightarrow{x_{g}z}$, then $\overrightarrow{x_{g}z}=\overrightarrow{x_{g}y_{g}}$ and $w_{g}(y_{g+1})=y_{g}$. Hence $\phi(\underline{x})=\underline{y}=(y_{g})\in X$.
It is left to prove that $\phi$ is an homeomorphism. For the injectivity, take $\underline{x}=(x_{g}),\underline{\tilde{x}}=(\tilde{x}_{g})\in X$ with $\underline{x}\neq\underline{\tilde{x}}$. We start proving that if $\underline{x}\neq\underline{q}$ then $\phi(\underline{x})\neq\underline{q}$. If not, $\phi(\underline{x})=\underline{y}=(y_{g})=\underline{q}=(q_{g})$ and for all $g$, $y_{g}=q_{g}$. Since $\underline{x}\neq\underline{q}$, there exists $g_{0}$ such that, for all $g\geq g_{0}$ $x_{g}\neq q_{g}$. In particular, for all $g\geq g_{0}$, $x_{g}$ either is the precursor of $q_{g}$ in the $c$-loop or $x_{g}$ is the precursor of $q_{g}$ in the $h$-loop. Denote this loop by $\lambda_{g}$. Take $g\geq g_{0}+1$, then there exists $z_{g}\in\lambda_{g}\setminus q_{g}$ such that $\overrightarrow{z_{g}q_{g}}\subset w_{g-1}^{-1}(q_{g-1})\cap\lambda_{g}$. Observe that $x_{g}$ cannot be contained in $\overrightarrow{z_{g}q_{g}}$, otherwise $w_{g-1}(x_{g})=q_{g-1}=x_{g-1}$. This contradicts the fact that $x_{g-1}\neq q_{g-1}$. Hence, if $\underline{x}\neq\underline{q}$, then $\phi(\underline{x})\neq\underline{q}$.
@@
In order to prove that $\phi$ is a surjection, let $\underline{y}=(y_{g})\in X$. If $\underline{y}=\underline{q}$, then $\phi(\underline{q})=\underline{y}$. Suppose $\underline{y}\neq\underline{q}$. By Lemma 26, there exists $g_{0}$ such that, for all $g\geq g_{0}$, $y_{g}\neq q_{g}$ and $y_{g}$ has a unique precursor $x_{g}$. If $w_{g}(\overrightarrow{x_{g+1}y_{g+1}})=\overrightarrow{z_{g}y_{g}}$, by the uniqueness of the predecessor $z_{g}=x_{g}$ and $w_{g}(x_{g+1})=x_{g}$. Hence, the sequence $\underline{x}=(x_{g})\in X$ and $\phi(\underline{x})=\underline{y}$.
@@
Case 1. If $\underline{x}\neq\underline{q}$ then, by Lemma 26 there exists $g_{0}$ such that for all $g\geq g_{0}$ $x_{g}\neq q_{g}$ and $\phi(x_{g})=y_{g}$ is the unique successor of $x_{g}$. Moreover, since $\underline{x}_{n}\to\underline{x}$, there exists $n_{0}$ such that, for all $n\geq n_{0}$, $x_{g,n}=x_{g}$. This says that for all $n\geq n_{0}$, $y_{g}$ is also the successor of $x_{n,g}$ and $\phi(x_{g,n})=y_{g,n}=y_{g}$. Hence, $\phi(\underline{x}_{n})\to\underline{y}=\phi(\underline{x})$.
Case 2. If $\underline{x}=\underline{q}$, then for all $g$, $x_{g}=q_{g}$. Choose $g$. Since $\underline{x}_{n}\to\underline{x}$, there exists $n_{0}$ such that, for all $n\geq n_{0}$, $x_{g+1,n}=q_{g+1}=x_{g+1}$. Let $y_{g+1,n}$ be the successor of $x_{g+1,n}=q_{g+1}$. Then $y_{g+1,n}\in Q_{g+1}\bigcup C_{g+1}(1)\bigcup H_{g+1}(1)\subset w_{g}^{-1}(q_{g})$. This implies that $y_{g,n}=w_{g+1}(y_{g+1,n})=q_{g}$. Hence, for all $n\geq n_{0}$, $y_{g,n}=q_{g}$ and $\phi(x_{g,n})=q_{g}$. In conlcusion $\phi(\underline{x})\to\underline{q}$. The construction shows that the homeomorphism is unique. ∎
@@ 10.2 Diagram
In this section we build a diagram structure as in Figure 12. This is the the core of the proof of Theorem 4.
@@
Orserve that, for all $g$, $\pi_{g+1}\circ w_{g}=i_{g}\circ\pi_{g}$. Moreover, for all $g$, let $j_{g}:A\to\mathcal{D}_{g}$ be the inclusion and let $j:A\to\mathcal{D}$. By construction the map $j$ is an homeomorphism. Similarly define the projection $\pi:\mathcal{D}\to X$.
Lemma 27.
@@
Let $\underline{x}=(x_{g})\in X$ and for all $g$, let $\pi^{-1}_{g}(x_{g})=U_{g}\subset\mathcal{D}_{g}$. Since for all $g$, $\pi_{g+1}\circ w_{g}=i_{g}\circ\pi_{g}$, it follows that $\dots\subset U_{g+1}\subset U_{g}$. Let ${a}\in\bigcap U_{g}$. Then $(a)\in\mathcal{D}$ and $\pi((a))=\underline{x}$. This proves that $\pi$ is a surjective map.
Moreover, since $\text{mesh}(U_{g})$ tends to zero, then $\left\{a\right\}=\bigcap U_{g}$. Hence $\pi^{-1}(\underline{x})=(a)$.
The proof is complete. ∎
@@
Let $a\in A$. If $a=p$, then $h(f((p))=h(p)=q$ and $\phi(h(p))=\phi(q)=q$.
Assume $a\neq p$, $a=\bigcap_{g}U_{g}$ where for all $g$, $U_{g}$ is a set in $\mathcal{D}_{g}$. Let $x_{g}=\pi_{g}(U_{g})$. By definition $h(a)=\underline{x}=(x_{g})$. Since $a\neq p$, there exists $g_{0}$ such that, for all $g\geq g_{0}$, $U_{g}\neq Q_{g}$ and $x_{g}\neq q_{g}$. In particular, $x_{g}$ has a unique successor $y_{g}$. Let $V_{g}=\pi^{-1}(y_{g})$. Since $\overrightarrow{x_{g}y_{g}}$ is an edge, $f(U_{g})\subset V_{g}$. Hence $\phi(h(a))=\phi(\underline{x})=(y_{g})=f(h(a))$. ∎
@@
Theorem 4.
If $f$ and $\tilde{f}$ are in the same leave $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets.
@@
The dynamical partitions of both maps $f$ and $\tilde{f}$ give rise to the same $(1,N_{g},M_{g})$-directed graph and graph morphism $w_{g}:X_{g+1}\mapsto X_{g}$, for all $g\geq 1$. According to Lemma 28 both $A(f)$ and A(f~)A($\tilde${f}) are conjugate to the corresponding $X$. ∎
Remark 15.
Probably, the Cantor sets of maps on different leaves are not conjugated.
@@ 10.3 Codimension-1 Stability
In this section we prove that the leaves of the lamination constructed above are the loci of infinitely renormalizable maps in the following sense.
Recall, from the proof of Lemma 13, there exists a smooth function
$m_{1}:S_{g}\mapsto\mathbb{R}^{2}$
@@
Definition 12.
Let $\ell$ be a leave of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$. A map $f\in S_{g}$ is renormalizable of type $(1,N_{g+1},M_{g+1})$ if $m_{7}(f)\in H_{g}(0)(f)$. Moreover, $f$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ if it is renormalizable of type $(1,N_{g+1},M_{g+1})$ for all $g\geq 1$.
Definition 13.
Let $f$ be an infinitely renormalizable map of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$. The Cantor set $A(f)$ associated to $f$ is codimension-$k$ stable if there exists a codimension-$k$ manifold $W$ in parameter space containing $f$ such that every map in $W$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ and there does not exists a manifold of codimension k1k-1 with the same property. The manifold $W$ is called the manifold of stability.
The following theorem proves Theorem A-(vii).
Theorem 5.
Let $\ell$ be a leave of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$. Every map $f\in\ell$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ and $A(f)$ is codimension-$1$ stable. Moreover, $\ell$ is the manifold of stability.
@@ 11 Invariant Measures
In this section we prove that each set $A$ has a unique invariant probability measure. In particular, the following theorem proves Theorem A-(iv).
Theorem 6.
@@
The following corollary proves Theorem A-(i).
Corollary 1.
@@
If $A$ is hyperbolic, then $A$ is equivalent to a subshift of finite type. Since $A$ has a dense orbit, this subset of finite type, is irreducible, i.e. it has a continuum of ergodic invariant measures. This contradicts Theorem 6. (Another reason is that $A$ has only one periodic orbit.) ∎
The proof of Theorem 6 requires some preparation. Let $X$ be the $(1,N,M)$ directed graph. We denote by $\delta^{q}$ the measure on $X$ which assigns unit mass to the vertex of the $q$-loop, $\lambda^{q}$. Similarly, we denote by $\delta^{c}$ the measure on $X$ which assigns unit mass to each vertex of the $c$-loop, $\lambda^{c}$ and by $\delta^{h}$ the measure which assigns unit mass to each vertex of the $h$-loop, $\lambda^{h}$. In follows that $\delta^{q}$ is a probability measure, $\delta^{c}(X)=N$ and $\delta^{h}(X)=M$.
@@
As in Remark 14, the winding of $w_{g}$ is given by
$\underline{w}\left(w_{g}\right)=\left(q,{q}^{\underline{\ell}_{g}}{c}{q}^{\theta_{g}n_{g}}{c}{q}^{n_{g}}{c}{q}^{n_{0,g}}{h}{q}^{\overline{\ell}_{g}},{q}^{\underline{\ell}_{g}}{h}{q}^{\overline{\ell}_{g}}\right).$
Then the matrix for $w_{g*}$, in the standard basis of $\mathfrak{M}_{g}$ and $\mathfrak{M}_{g+1}$ is
$W_{g}=\left(\begin{matrix}1&\underline{\ell}_{g}+\theta_{g}n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}&\underline{\ell}_{g}+\overline{\ell}_{g}\\ 0&3&0\\ 0&1&1\end{matrix}\right)$
@@
Let $\mathcal{I}$ be the set of invariant probability measures of $f:X\to X$. If $\mu\in\mathcal{I}$ denote the restriction of the measure to the $\varsigma$-algebra of $X_{g}$ by
@@
$\mathcal{I}\ni\mu\mapsto(\mu_{g})\in\mathcal{P}$
is a bijection. For more details on this construction see LABEL:GM.
@@
$\underline{\delta}^{c}_{g}=\frac{1}{N_{g}},\underline{\delta}^{h}_{g}=\frac{1}{M_{g}},\underline{\delta}^{q}_{g}=\delta^{q}_{g}.$
@@
Define the norm $|.|$ on $\mathfrak{M}_{g}$ by
$\mu=\max\{|x|,|y|,|z|\},$
@@
The map $W_{g}:\mathcal{P}_{g+1}\to\mathcal{P}_{g}.$ contracts towards $\underline{\delta}^{q}_{g}$. Namely,
Lemma 29.
@@
Observe, $W_{g}\underline{\delta}^{q}_{g+1}=\underline{\delta}^{q}_{g}$. Hence, it suffices to prove the estimate for $\mu=\underline{\delta}^{c}_{g}$ and $\mu=\underline{\delta}^{h}_{g}$. We will only present the estimate for $\mu=\underline{\delta}^{c}_{g}$, the other is similar.
Observe,
$W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{q}_{g})=\frac{\underline{\ell}_{g}+\theta_{g}n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}}{N_{g+1}}$
@@
$N_{g+1}=\underline{\ell}_{g}+N_{g}+\theta_{g}n_{g}+N_{g}+n_{g}+N_{g}+n_{0,g}+M_{g}+\overline{\ell}_{g}$
@@
Proof of Theorem 6: Let $\mu\in\mathcal{P}$, say $\mu=(\mu_{g})$. Recall, this measure corresponds to an invariant probability of $f$. We will show that $\mu_{g}=\underline{\delta}^{q}_{g}$, for all $g\geq 1$. The previous lemma implies that for all $s\geq 1$
@@
Hence, $\mu_{g}=\underline{\delta}^{q}_{g}$, for all $g\geq 1$. This means that the only invariant measure of $f:A\to A$ is the unit mass on its fixed point.
12 Properties of the Lamination: Collet-Eckmann and Newhouse points
In this section, we prove the existence of a sub-lamination whose leaves have an interesting property. In particular we prove Theorem B and Theorem C.
The invariant non-hyperbolic Cantor set, associated to each map in the leaves of this sublamination have a Collet-Eckmann point, which is dense in the Cantor set. Let us start by giving the formal definition of a Collet-Eckmann point.
Definition 14.

### v3

@@ 9.3 Properties of the set AA
$\bigcap_{g}\left(Q_{g}\cup\bigcup_{j=1}^{M_{g}-1}H_{g}^{j}\right)=O_{2}.$
@@
Proposition 9.3.6.
Let $\ell$ be a leaf of the lamination $\mathfrak{L}$ and let $f\in\ell$. Then the map $f:A(f)\mapsto A(f)$ is an homeomorphism.
@@
Since $f$ is injective, its restriction to $A$ is injective as well. We only need to prove that $f(A(f))=A(f)$. Observe that for all $x$, $f(\omega(x))=\omega(x)$. In particular, when $x=c$, $f(\omega(c))=\omega(c)$ and by Lemma 9.3.4 we obtain $f(A(f))=A(f)$. The proof is complete. ∎
@@ 10 Topological Equivalence along the Leaves
The aim of this section is to prove that two maps which are in the same leaf of the lamination are conjugate on their respective Cantor sets, see Theorem 10.2.3. In order to do that, we need to build a topological model for the invariant Cantor sets.
@@ 10.1 Inverse limits
Take a leaf $\ell$ of the lamination $\mathfrak{L}$ as constructed in (9.1.1) and (9.1.2). By the construction following Proposition 7.4.1 we know that for all $g$ there is a $(1,N_{g},M_{g})$-directed graph and a graph morphism $w_{g}:X_{g+1}\mapsto X_{g}$. We can then define the inverse limit as
@@
with the corresponding topology. A sequence $\left\{\underline{x}_{n}\right\}_{n\in\mathbb{N}}=\left\{\left(x_{g,n}\right)\right\}_{n\in\mathbb{N}}\subset X$ converges to a point $\underline{x}=\left(x_{g}\right)\in X$, namely
@@
In particular $X$ is a Cantor set.
Lemma 10.1.1.
Let $\underline{x}\in X$, $\underline{x}=(x_{g})$. If there exists $g_{0}$ such that $x_{g_{0}}=q_{g_{0}}$ then for all $g\leq g_{0}$, $x_{g}=q_{g}$.
@@
In the following proposition we prove the existence of an homeomorphism on the inverse limit defined by the directed graphs.
Proposition 10.1.2.
@@
Let $\underline{x}=(x_{g})\in X$. By Lemma 10.1.1, either $\underline{x}=\underline{q}=(q_{g})$ or there exists $g_{0}$ such that, for all $g\geq g_{0}$, $x_{g}\neq q_{g}$. For the case when $\underline{x}=\underline{q}$, define $\phi(\underline{x})=\underline{q}$. In the other case define $\phi(\underline{x})=\underline{y}=(y_{g})$ where $y_{g}=q_{g}$ for all $g<g_{0}$ and $y_{g}$ is the vertex in the unique outgoing edge starting in $x_{g}$. Since $w_{g}(x_{g+1})=x_{g}$ and since $x_{g}$ has a unique outgoing edge, if $w_{g}(\overrightarrow{x_{g+1}y_{g+1}})=\overrightarrow{x_{g}z}$, then $\overrightarrow{x_{g}z}=\overrightarrow{x_{g}y_{g}}$ and $w_{g}(y_{g+1})=y_{g}$. Hence $\phi(\underline{x})=\underline{y}=(y_{g})\in X$.
It is left to prove that $\phi$ is an homeomorphism. We start proving that if $\underline{x}\neq\underline{q}$ then $\phi(\underline{x})\neq\underline{q}$. If not, $\phi(\underline{x})=\underline{y}=(y_{g})=\underline{q}=(q_{g})$ and for all $g$, $y_{g}=q_{g}$. Since $\underline{x}\neq\underline{q}$, there exists $g_{0}$ such that, for all $g\geq g_{0}$ $x_{g}\neq q_{g}$. In particular, for all $g\geq g_{0}$, $x_{g}$ either is the predecessor of $q_{g}$ in the $c$-loop or $x_{g}$ is the predecessor of $q_{g}$ in the $h$-loop. Denote this loop by $\lambda_{g}$. Take $g\geq g_{0}+1$, then there exists $z_{g}\in\lambda_{g}\setminus q_{g}$ such that $\overrightarrow{z_{g}q_{g}}\subset w_{g-1}^{-1}(q_{g-1})\cap\lambda_{g}$. In particular we can choose $z_{g}$ with the property that the length of $\overrightarrow{z_{g}q_{g}}$ is $\underline{\ell}_{g}$ which is larger than $2$. Hence, the predecessor of $q_{g}$ in $\lambda_{g}$ is contained in this path, i.e. $x_{g}$ is in $\overrightarrow{z_{g}q_{g}}$. This implies that $w_{g-1}(x_{g})=q_{g-1}=x_{g-1}$. Contradiction. We proved that if $\underline{x}\neq\underline{q}$, then $\phi(\underline{x})\neq\underline{q}$.
@@
In order to prove that $\phi$ is a surjection, let $\underline{y}=(y_{g})\in X$. If $\underline{y}=\underline{q}$, then $\phi(\underline{q})=\underline{y}$. Suppose $\underline{y}\neq\underline{q}$. By Lemma 10.1.1, there exists $g_{0}$ such that, for all $g\geq g_{0}$, $y_{g}\neq q_{g}$ and $y_{g}$ has a unique predecessor $x_{g}$. Let $w_{g}(\overrightarrow{x_{g+1}y_{g+1}})=\overrightarrow{z_{g}y_{g}}$. Then, by the uniqueness of the predecessor, $z_{g}=x_{g}$ and $w_{g}(x_{g+1})=x_{g}$. Hence, the sequence $\underline{x}=(x_{g})\in X$ and $\phi(\underline{x})=\underline{y}$.
@@
Case 1. If $\underline{x}\neq\underline{q}$ then, by Lemma 10.1.1 there exists $g_{0}$ such that for all $g\geq g_{0}$ $x_{g}\neq q_{g}$ and $\phi(x_{g})=y_{g}$ is the unique successor of $x_{g}$. Moreover, since $\underline{x}_{n}\to\underline{x}$, there exists $n_{0}$ such that, for all $n\geq n_{0}$, $x_{g,n}=x_{g}$. This says that for all $n\geq n_{0}$, $y_{g}$ is also the successor of $x_{n,g}$ and $\phi(x_{g,n})=y_{g,n}=y_{g}$. Hence, $\phi(\underline{x}_{n})\to\underline{y}=\phi(\underline{x})$.
Case 2. If $\underline{x}=\underline{q}$, then for all $g$, $x_{g}=q_{g}$. Choose $g$. Since $\underline{x}_{n}\to\underline{x}$, there exists $n_{0}$ such that, for all $n\geq n_{0}$, $x_{g+1,n}=q_{g+1}=x_{g+1}$. Let $y_{g+1,n}$ be the successor of $x_{g+1,n}=q_{g+1}$. Then $y_{g+1,n}\in Q_{g+1}\bigcup C_{g+1}^{1}\bigcup H_{g+1}^{1}\subset w_{g}^{-1}(q_{g})$. This implies that $y_{g,n}=w_{g+1}(y_{g+1,n})=q_{g}$. Hence, for all $n\geq n_{0}$, $y_{g,n}=q_{g}$. In conlcusion $\phi(\underline{x}_{n})\to\underline{q}$.
The construction shows that the homeomorphism is unique. ∎
@@ 10.2 Diagram
In this section we build a diagram as in Figure 13. This is the the core of the proof of Theorem 10.2.3.
@@
Observe that, for all $g$, $\pi_{g+1}\circ w_{g}=i_{g}\circ\pi_{g}$. Moreover, for all $g$, let $j_{g}:A\to\mathcal{D}_{g}$ be the inclusion and denote the corresponding limit by $j:A\to\mathcal{D}$. By construction the map $j$ is an homeomorphism. Similarly define the projection $\pi:\mathcal{D}\to X$.
Lemma 10.2.1.
@@
Let $\underline{x}=(x_{g})\in X$ and for all $g$, let $\pi^{-1}_{g}(x_{g})=U_{g}\subset\mathcal{D}_{g}$. Since for all $g$, $\pi_{g+1}\circ w_{g}=i_{g}\circ\pi_{g}$, it follows that $\dots\subset U_{g+1}\subset U_{g}$. Let ${a}\in\bigcap U_{g}$. Then $a\in\mathcal{D}$ and $\pi(a)=\underline{x}$. This proves that $\pi$ is a surjective map. Moreover, since $\text{mesh}(U_{g})$ tends to zero, then $\left\{a\right\}=\bigcap U_{g}$. Hence $\pi^{-1}(\underline{x})=(a)$. The proof is complete. ∎
@@
Let $a\in A$. If $a=p$, then $h(f(p))=h(p)=q$ and $\phi(h(p))=\phi(q)=q$. Assume $a\neq p$, $a=\bigcap_{g}U_{g}$ where for all $g$, $U_{g}$ is a set in $\mathcal{D}_{g}$. Let $x_{g}=\pi_{g}(U_{g})$. By definition $h(a)=\underline{x}=(x_{g})$. Since $a\neq p$, there exists $g_{0}$ such that, for all $g\geq g_{0}$, $U_{g}\neq Q_{g}$ and $x_{g}\neq q_{g}$. In particular, $x_{g}$ has a unique successor $y_{g}$. Let $V_{g}=\pi^{-1}(y_{g})$. Since $\overrightarrow{x_{g}y_{g}}$ is an edge, $f(U_{g})\subset V_{g}$. Hence $\phi(h(a))=\phi(\underline{x})=(y_{g})=f(h(a))$. ∎
@@
Theorem 10.2.3.
If $f$ and $\tilde{f}$ are in the same leaf $\ell$ of the lamination $\mathfrak{L}$ then they are conjugate on their respective Cantor sets.
@@
The dynamical partitions of both maps $f$ and $\tilde{f}$ give rise to the same $(1,N_{g},M_{g})$-directed graph and graph morphism $w_{g}:X_{g+1}\mapsto X_{g}$, for all $g\geq 1$. According to Lemma 10.2.2 both $A(f)$ and A(f~)A($\tilde${f}) are conjugate to the corresponding $X$. ∎
Remark 10.2.4.
Probably, the Cantor set of maps on different leaves are not conjugated.
@@ 10.3 Codimension-1 Stability
In this section, following the Remark 7.3.2, we prove that the leaves of the lamination constructed above are the loci of infinitely renormalizable maps in the following sense. Let $\ell=\cap S_{g}$ be a leaf of $\mathfrak{L}$. Recall, from the proof of Lemma 5.1.1, there exists a smooth function
$m_{1}:S_{g}\to\mathbb{R}^{2},$
@@
Definition 10.3.1.
Let $\ell=\cap S_{g}$ be a leaf of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$. A map $f\in S_{g}$ is renormalizable of type $(1,N_{g+1},M_{g+1})$ if $m_{7}(f)\in H_{g}^{0}(f)$. Moreover, $f$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ if it is renormalizable of type $(1,N_{g+1},M_{g+1})$ for all $g\geq 1$.
Definition 10.3.2.
Let $f$ be an infinitely renormalizable map of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$. The Cantor set $A(f)$ associated to $f$ is codimension-$k$ stable if there exists a codimension-$k$ manifold $W$ in parameter space containing $f$ such that every map in $W$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ and there does not exists a manifold of codimension-k1k-1 with the same property. The manifold $W$ is called the manifold of stability.
The following theorem confirms Theorem A-(vii).
Theorem 10.3.3.
Let $\ell$ be a leaf of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$. Every map $f\in\ell$ is infinitely renormalizable of type $\left\{\left(1,N_{g},M_{g}\right)\right\}_{g=1}^{\infty}$ and $A(f)$ is codimension-$1$ stable. Moreover, $\ell$ is the manifold of stability.
@@ 11 Invariant Measures
In this section we prove that each set $A$ has a unique invariant probability measure. In particular, the following theorem confirms Theorem A-(iv).
Theorem 11.0.1.
@@
Corollary 11.0.2.
@@
If $A$ is hyperbolic, then $A$ is equivalent to a subshift of finite type. Since $A$ has a dense orbit, this subset of finite type, is irreducible, i.e. it has a continuum of ergodic invariant measures. This contradicts Theorem 11.0.1. (Another reason is that $A$ has only one periodic orbit.) ∎
The proof of Theorem 11.0.1 requires some preparation. Let $X$ be the $(1,N,M)$ directed graph. We denote by $\delta^{q}$ the measure on $X$ which assigns unit mass to the vertex of the $q$-loop, $\lambda^{q}$. Similarly, we denote by $\delta^{c}$ the measure on $X$ which assigns unit mass to each vertex of the $c$-loop, $\lambda^{c}$ and by $\delta^{h}$ the measure which assigns unit mass to each vertex of the $h$-loop, $\lambda^{h}$. In follows that $\delta^{q}$ is a probability measure, $\delta^{c}(X)=N$ and $\delta^{h}(X)=M$.
@@
As in Remark 7.4.4, the winding of $w_{g}$ is given by
$\underline{\omega}\left(w_{g}\right)=\left(q,{q}^{\underline{\ell}_{g}}{c}{q}^{\theta n_{g}}{c}{q}^{n_{g}}{c}{q}^{n_{0,g}}{h}{q}^{\overline{\ell}_{g}},{q}^{\underline{\ell}_{g}}{h}{q}^{\overline{\ell}_{g}}\right).$
Then the matrix for $w_{g*}$, in the standard basis of $\mathfrak{M}_{g}$ and $\mathfrak{M}_{g+1}$ can be easily read off from $\underline{\omega}\left(w_{g}\right)$ and it is
$W_{g}=\left(\begin{matrix}1&\underline{\ell}_{g}+\theta n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}&\underline{\ell}_{g}+\overline{\ell}_{g}\\ 0&3&0\\ 0&1&1\end{matrix}\right).$
@@
Let $\mathcal{I}$ be the set of invariant probability measures of $f:X\to X$. If $\mu\in\mathcal{I}$ then the restriction of the measure to the $\sigma$-algebra of $X_{g}$ is denoted by
@@
$\mathcal{I}\ni\mu\mapsto(\mu_{g})\in\mathcal{P}$ (11.0.3)
is a bijection. For more details on this construction see [32].
@@
$\underline{\delta}^{c}_{g}=\frac{1}{N_{g}}{\delta}^{c}_{g},\text{ }\underline{\delta}^{h}_{g}=\frac{1}{M_{g}}{\delta}^{h}_{g},\text{ }\underline{\delta}^{q}_{g}=\delta^{q}_{g}.$
@@
Define the norm $|\mu|$ of $\mu\in\mathfrak{M}_{g}$ by
$|\mu|=\max\{|x|,|y|,|z|\},$
@@
The map $W_{g}:\mathcal{P}_{g+1}\to\mathcal{P}_{g}$ contracts towards $\underline{\delta}^{q}_{g}$. Namely,
Lemma 11.0.4.
@@
Observe, $W_{g}\underline{\delta}^{q}_{g+1}=\underline{\delta}^{q}_{g}$. Hence, it suffices to prove the estimate for $\mu=\underline{\delta}^{c}_{g}$ and $\mu=\underline{\delta}^{h}_{g}$. We will only present the estimate for $\mu=\underline{\delta}^{c}_{g}$, the other is similar. Observe,
$W_{g}\underline{\delta}^{c}_{g+1}(\lambda^{q}_{g})=\frac{\underline{\ell}_{g}+\theta n_{g}+n_{g}+n_{0,g}+\overline{\ell}_{g}}{N_{g+1}}.$
@@
$N_{g+1}=\underline{\ell}_{g}+N_{g}+\theta n_{g}+N_{g}+n_{g}+N_{g}+n_{0,g}+M_{g}+\overline{\ell}_{g}$
@@
Proof of Theorem 11.0.1: Observe that the conditions in Lemma 11.0.4 are guaranteed by Remark 9.1.4. Let $\mu\in\mathcal{I}$ be an invariant probability measure of $f$. In particular, $\mu$ corresponds to $(\mu_{g})\in\mathcal{P}$, see (11.0.3). We will show that $\mu_{g}=\underline{\delta}^{q}_{g}$, for all $g\geq 1$. The previous lemma implies that for all $s\geq 1$
@@
Hence, $\mu_{g}=\underline{\delta}^{q}_{g}$, for all $g\geq 1$. This means that the only invariant measure of $f:A\to A$ is the unit mass on its fixed point. ∎
12 Collet-Eckmann Lamination and Newhouse points
In the previous sections we observed many different phenomena. In this section we show that all these phenomena occur in one and the same lamination. In particular there exists a lamination with a stable non-hyperbolic Cantor set which contains a Collet-Eckmann point with a dense orbit. Moreover each leave has also a Newhouse point. At this Newhouse points there is an intriguing coexistence of stable and unstable dynamics. This section confirms Theorem B and Theorem C.
Let us start by giving the formal definition of a Collet-Eckmann point.
Definition 12.0.1.